# Equivalence principles in Weyl transverse gravity

### v1

@@ Abstract
There exist two consistent theories of massless, self-interacting gravitons, which differ by their local symmetries: general relativity and Weyl transverse gravity. We show that these two theories are also the only two metric descriptions of gravity in $4$ spacetime dimensions which obey the equivalence principle for test gravitational physics. We further analyse how the weaker formulations of the equivalence principle are realised in Weyl transverse gravity (and its generalisations). The analysis sheds light on the behaviour of matter fields in this theory.
Equivalence principle, Weyl transverse gravity, unimodular gravity
@@ I Introduction
The equivalence principle (EP) historically played an important role in guiding the development of general relativity [1]. It essentially states that suitably defined “test physics” is not affected by the presence of a background gravitational field [2, 3, 4]. In this context, test physics refers to localised physical processes whose effect on the background gravitational field is negligible. There exists a hierarchy of different formulations of the EP, depending on which processes one chooses as admissible test physics. For instance, the weak EP only concerns geodesic motion of test particles with negligible self-gravity. The most restrictive version known as the strong EP extends to arbitrary test physics including gravitational phenomena, such as the propagation of gravitational waves (provided that their backreaction on the background spacetime can be neglected). The classification of the various formulations of the EP we adopt follows Ref. [2]. In general, we can separate the EPs into two groups. The ones concerning non-gravitational test physics constrain the interaction between matter and fixed gravitational background, essentially determining the kinematics of gravity [3, 4]. The EPs which extend to gravitational test physics then also impose constraints on gravitational dynamics compatible with them [5].
@@
Nowadays, the interest in the EP lies in the experimental and observational studies of its validity [6, 7, 8] and in studies of its validity in quantum physics [9, 10, 11, 12]. Moreover, the various candidate theories of gravity can be classified according to which formulations of the EP they are compatible with [5, 2]. In particular, the only theories known to obey the strong EP are the Lanczos-Lovelock theories of gravity [13, 14, 15]. In $4$ spacetime dimensions, general relativity is the single theory that has been previously shown to incorporate the strong EP.
@@
Herein, we show that the strong EP is compatible with at least one other gravitational theory in $4$ dimensions, namely, with Weyl transverse gravity [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. This theory has the same classical solutions as general relativity. However, its equations of motion are traceless and the theory has different local symmetries. Rather than being invariant under arbitrary diffeomorphisms (Diff-invariant), it is instead invariant under the subgroup of spacetime volume preserving diffeomorphisms and under Weyl transformations (WTDiff-invariant). Both theories thus have $D\left(D+1\right)/2$ local symmetries, with $D$ being the spacetime dimension. Weyl transverse gravity first emerged from the construction of a theory of massless, self-interacting gravitons [18, 21, 23, 24, 26]. There exist precisely two such theories that do not involve any gauge fixing and their only propagating degrees of freedom are those of the graviton. These theories differ by the choice of the symmetry group, with Diff invariance leading to general relativity and WTDiff invariance to Weyl transverse gravity.
@@
Weyl transverse gravity passes all the experimental and observational tests up to date (having the same classical solutions as general relativity), provides a consistent theory for self-interacting gravitons, and offers a robust solution for one of the problems related to the value of the cosmological constant [16, 22, 24]. Thence, it represents a viable and relevant alternative to general relativity. By showing that Weyl transverse gravity satisfies the strong EP, we offer a further argument for singling it out as a competitor of general relativity.
@@
In addition, Weyl transverse gravity is actually favoured over general relativity by approaches that derive the gravitational dynamics from thermodynamic considerations [29] (for a broader context on such thermodynamic derivations see, e.g. [30, 31, 32]). Since one of the key assumptions of any local thermodynamic derivation is the validity of the strong EP [33], checking that this assumption is consistent with Weyl transverse gravity provides an additional motivation for this work.
@@
Aside from studying the strong EP, we also discuss various weaker formulations of the EP and show how they are incorporated in Weyl transverse gravity. Furthermore, we study their validity even in more general theories of gravity with WTDiff symmetry.
@@
The paper is organised as follows. In section II, we review the basics of Weyl transverse gravity and of more general WTDiff-invariant theories. Section III discusses the EPs restricted to non-gravitational test physics. We then consider the EPs for gravitational test physics in section IV. Finally, section V summarises our findings and discusses possible future development.
@@
We work in an arbitrary spacetime dimension $D$ (unless specified otherwise) with a metric signature (,+,,+)(-,+,...,+)( - , + , … , + ). We set $c=k_{\text{B}}=G=\hbar=1$. Lowercase Greek letters are used for abstract spacetime indices. Other conventions follow [34].
II Weyl transverse gravity and its generalisations
@@
Herein, we recall the main features of Weyl transverse gravity and of WTDiff-invariant theories of gravity in general. For a detailed review, we refer the readers to [26].
@@
where $\mathfrak{g}$ denotes the metric determinant. Both $\sqrt{-\mathfrak{g}}$ and $\omega$ are scalar densities of weight $+1$, ensuring that $\tilde{g}_{\mu\nu}$ (which depends on their ratio) is a tensor. This auxiliary metric can be understood as a restriction of $g_{\mu\nu}$ to the unimodular gauge, $\sqrt{-\mathfrak{g}}=\omega$. We stress that we treat $\tilde{g}_{\mu\nu}$ as a mere notational device, keeping $g_{\mu\nu}$ as the dynamical field. To ensure we always work with WTDiff-invariant expressions, raising and lowering of indices is performed with $\tilde{g}_{\mu\nu}$ and its inverse metric $\tilde{g}^{\mu\nu}$.
@@
~$\tilde{\nabla}_{\rho}g_{\mu\nu}=g_{\mu\nu}\frac{2}{D}\nabla_{\rho}\ln\frac{\sqrt{-\mathfrak{g}}}{\omega}.$ (3)
Since the result is a tensor product of the metric and a gradient of a scalar function, it follows that $\tilde{\Gamma}^{\mu}_{\;\>\nu\rho}$ is an integrable Weyl connection for the dynamical metric, as well as for any metric conformally related to it [35, 36]. This observation will play a role in our discussion of the EPs. Using the Weyl connection (2), we introduce an auxiliary, WTDiff-invariant Riemann tensor
@@
The simplest action one can construct from the auxiliary metric and the corresponding Riemann tensor is that of Weyl transverse gravity, i.e.,
@@
Since the spacetime volume measure is non-dynamical, adding any constant term to the Weyl transverse gravity Lagrangian amounts simply to shifting the action by a constant and does not affect dynamics in any way. Hence, we are free to set this constant term to zero in the following. This marks a departure from general relativity, where a constant term in the Lagrangian corresponds to the cosmological constant.
@@
where $\sigma$ is an arbitrary scalar function. The volume measure $\omega$ and the matter fields are by definition unaffected by Weyl transformations, ensuring the Weyl invariance of $\tilde{g}_{\mu\nu}$. Furthermore, $I_{\text{WTG}}$ is invariant under transverse diffeomorphisms, but not under longitudinal ones. However, we must be careful to consider the appropriate notion of transversality. The usual condition on the generator $\xi^{\mu}$ of transverse diffeomorphisms, $\nabla_{\mu}\xi^{\mu}=0$, is not Weyl invariant. Thus, it cannot be satisfied in every Weyl frame simultaneously, making it unsuitable for Weyl transverse gravity. Instead, one must define transversality with respect to the Weyl invariant covariant derivative. Hence, the appropriate transversality condition reads
~$\tilde{\nabla}_{\mu}\xi^{\mu}=0\qquad\iff\qquad\nabla_{\mu}\xi^{\mu}=\xi^{\mu}\partial_{\mu}\ln\frac{\sqrt{-\mathfrak{g}}}{\omega}.$ (7)
Since the Lie derivative of the volume $D$-form $\bm{\omega}$ yields £$\pounds_{\xi}\bm{\omega}=\bm{\omega}\tilde{\nabla}_{\mu}\xi^{\mu}$ (this result can be obtained by direct computation and also follows from the fact that ~$\tilde{\nabla}_{\mu}\bm{\omega}=0$), we can understand the transversality condition as defining the volume preserving transformations. Transverse diffeomorphisms transform the dynamical metric $g_{\mu\nu}$ in the usual way
@@
To find equations of motion for Weyl transverse gravity, we vary action (5) with respect to the dynamical metric $g^{\mu\nu}$, obtaining traceless, WTDiff-invariant equations of motion
$\tilde{R}_{\mu\nu}-\frac{1}{D}\tilde{R}\tilde{g}_{\mu\nu}=8\pi G\left(\tilde{T}_{\mu\nu}-\frac{1}{D}\tilde{T}\tilde{g}_{\mu\nu}\right),$ (10)
where we define the WTDiff-invariant energy-momentum tensor
@@
~$\tilde{\nabla}_{\nu}\tilde{T}_{\mu}^{\;\>\nu}=\tilde{\nabla}_{\mu}\mathcal{J},$ (12)
@@
Bianchi identities allow us to restate the traceless equations of motion (10) in a divergenceless form reminiscent of the equations of motion of general relativity
@@
where $\Lambda$ is an arbitrary integration constant and the divergenceless tensor
@@
acts as the matter source for the Einstein tensor. Hence, $\tilde{T}^{\prime}_{\mu\nu}$ is the relevant tensor for coupling of matter fields to Weyl transverse gravity, as we will see in our discussion of the EP.
@@
The integration constant $\Lambda$ plays the role of the cosmological constant. In contrast to general relativity, $\Lambda$ is unrelated to any fixed parameter of the Lagrangian and is only defined on shell, generically taking different values for various solutions of the equations of motion. In other words, $\Lambda$ represents a global degree of freedom of the theory.
@@
In the same way that there exist modified Diff-invariant theories of gravity, there exist WTDiff-invariant theories generalising Weyl transverse gravity. Their Lagrangian can be any scalar constructed from the auxiliary metric, $\tilde{g}_{\mu\nu}$, the auxiliary Riemann tensor, $\tilde{R}^{\mu}_{\;\>\nu\rho\sigma}$, the Weyl covariant derivative ~$\tilde{\nabla}_{\mu}$ and some collection of matter fields, in principle non-minimally coupled. One can in fact show that for each local, WTDiff-invariant action there exists a corresponding local, Diff-invariant action which implies the same classical gravitational dynamics (except for the different behaviour of $\Lambda$) and vice versa [26]. In other words, there exist pairs of Diff and WTDiff-invariant theories of gravity that are mutually equivalent in the same sense in which general relativity is equivalent to Weyl transverse gravity. As a special case, there exist WTDiff-invariant Lanczos-Lovelock theories [13, 14] which are the only purely metric, WTDiff-invariant theories with second order equations of motion. They have the same classical solutions as the corresponding Diff-invariant Lanczos-Lovelock gravitational theories [15]. As we comment in section IV, Lanczos-Lovelock gravity, both Diff- and WTDiff-invariant, is also notable for obeying the strong equivalence principle.

### v2

# Equivalence principles in WTG

@@ Abstract
There exist two consistent theories of massless, self-interacting gravitons, which differ by their local symmetries: GR and WTG. We show that these two theories are also the only two metric descriptions of gravity in $4$ spacetime dimensions which obey the equivalence principle for test gravitational physics. We further analyse how the weaker formulations of the equivalence principle are realised in WTG (and its generalisations). The analysis sheds light on the behaviour of matter fields in this theory.
Equivalence principle, WTG, unimodular gravity
@@ I Introduction
The equivalence principle (EP) historically played an important role in guiding the development of GR [1]. It essentially states that suitably defined “test physics” is not affected by the presence of a background gravitational field [2, 3, 4]. In this context, test physics refers to localised physical processes whose effect on the background gravitational field is negligible. There exists a hierarchy of different formulations of the EP, depending on which processes one chooses as admissible test physics. For instance, the weak EP only concerns geodesic motion of test particles with negligible self-gravity. The most restrictive version known as the strong EP extends to arbitrary test physics including gravitational phenomena, such as the propagation of gravitational waves (provided that their backreaction on the background spacetime can be neglected). The classification of the various formulations of the EP we adopt follows Ref. [2]. In general, we can separate the EPs into two groups. The ones concerning non-gravitational test physics constrain the interaction between matter and fixed gravitational background, essentially determining the kinematics of gravity [3, 4]. The EPs which extend to gravitational test physics then also impose constraints on gravitational dynamics compatible with them [5].
@@
Nowadays, the interest in the EP lies in the experimental and observational studies of its validity [6, 7, 8] and in studies of its validity in quantum physics [9, 10, 11, 12]. Moreover, the various candidate theories of gravity can be classified according to which formulations of the EP they are compatible with [5, 2]. In particular, the only theory known to obey the strong EP is general relativity (GR).
@@
Herein, we show that the strong EP is compatible with at least one other gravitational theory in $4$ dimensions, namely, with Weyl transverse gravity (WTG) [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. This theory has the same classical solutions as GR [^1] Note here that explicitly WTG solutions are equivalent to the whole set of GR theories, one for each value of the cosmological constant, because in this theory the cosmological constant emerged as a constant of motion which can take different values while in GR the cosmological constant is fixed upon choosing its value in the action.. However, its equations of motion are traceless and the theory has different local symmetries. Rather than being invariant under arbitrary diffeomorphisms (Diff-invariant), it is instead invariant under the subgroup of spacetime volume preserving diffeomorphisms and under Weyl transformations (WTDiff-invariant). Both theories thus have $D\left(D+1\right)/2$ local symmetries, with $D$ being the spacetime dimension. WTG first emerged from the construction of a theory of massless, self-interacting gravitons [15, 18, 20, 21, 23]. There exist precisely two such theories that do not involve any gauge fixing and their only propagating degrees of freedom are those of the graviton. These theories differ by the choice of the symmetry group, with Diff invariance leading to GR and WTDiff invariance to WTG.
@@
WTG passes all the experimental and observational tests up to date (having the same classical solutions as GR), provides a consistent theory for self-interacting gravitons, and offers a robust solution for one of the problems related to the value of the cosmological constant [13, 19, 21]. Thence, it represents a viable and relevant alternative to GR. By showing that WTG satisfies the strong EP, we offer a further argument for singling it out as a competitor of GR.
@@
In addition, WTG is actually favoured over GR by approaches that derive the gravitational dynamics from thermodynamic considerations [26] (for a broader context on such thermodynamic derivations see, e.g. [27, 28, 29]). Since one of the key assumptions of any local thermodynamic derivation is the validity of the strong EP [30], checking that this assumption is consistent with WTG provides an additional motivation for this work.
@@
Aside from studying the strong EP, we also discuss various weaker formulations of the EP and show how they are incorporated in WTG. Furthermore, we study their validity even in more general theories of gravity with WTDiff symmetry.
@@
The paper is organised as follows. In section II, we review the basics of WTG and of more general WTDiff-invariant theories. Section III discusses the EPs restricted to non-gravitational test physics. We then consider the EPs for gravitational test physics in section IV. Finally, section V summarises our findings and discusses possible future development.
@@
We work in an arbitrary spacetime dimension $D$ (unless specified otherwise) with a metric signature (,+,,+)(-,+,...,+). We set $c=k_{\text{B}}=G=\hbar=1$. Lowercase Greek letters are used for abstract spacetime indices. Other conventions follow [31].
II WTG and its generalisations
@@
Herein, we recall the main features of WTG and of WTDiff-invariant theories of gravity in general. For a detailed review, we refer the readers to [23].
@@
where $\mathfrak{g}$ denotes the metric determinant. Both $\sqrt{-\mathfrak{g}}$ and $\omega$ are scalar densities of weight +1+1, ensuring that $\tilde{g}_{\mu\nu}$ (which depends on their ratio) is a tensor. This auxiliary metric can be understood as a restriction of $g_{\mu\nu}$ to the unimodular gauge, $\sqrt{-\mathfrak{g}}=\omega$. We stress that we treat $\tilde{g}_{\mu\nu}$ as a mere notational device, keeping $g_{\mu\nu}$ as the dynamical field. To ensure we always work with WTDiff-invariant expressions, raising and lowering of indices is performed with $\tilde{g}_{\mu\nu}$ and its inverse metric $\tilde{g}^{\mu\nu}$.
@@
$\tilde{\nabla}_{\rho}g_{\mu\nu}=g_{\mu\nu}\frac{2}{D}\nabla_{\rho}\ln\frac{\sqrt{-\mathfrak{g}}}{\omega}.$ (3)
Since the result is a tensor product of the metric and a gradient of a scalar function, it follows that $\tilde{\Gamma}^{\mu}_{\;\>\nu\rho}$ is an integrable Weyl connection for the dynamical metric, as well as for any metric conformally related to it (for details on Weyl connections see, e.g. [32]). This observation will play a role in our discussion of the EPs.
Using the Weyl connection (2), we introduce an auxiliary, WTDiff-invariant Riemann tensor
@@
The simplest action one can construct from the auxiliary metric and the corresponding Riemann tensor is that of WTG, i.e.,
@@
Since the spacetime volume measure is non-dynamical, adding any constant term to the WTG Lagrangian amounts simply to shifting the action by a constant and does not affect dynamics in any way. Hence, we are free to set this constant term to zero in the following. This marks a departure from GR, where a constant term in the Lagrangian corresponds to the cosmological constant.
@@
where $\sigma$ is an arbitrary scalar function. The volume measure $\omega$ and the matter fields are by definition unaffected by Weyl transformations, ensuring the Weyl invariance of $\tilde{g}_{\mu\nu}$. Furthermore, $I_{\text{WTG}}$ is invariant under transverse diffeomorphisms, but not under longitudinal ones. However, we must be careful to consider the appropriate notion of transversality. The usual condition on the generator $\xi^{\mu}$ of transverse diffeomorphisms, $\nabla_{\mu}\xi^{\mu}=0$, is not Weyl invariant. Thus, it cannot be satisfied in every Weyl frame simultaneously, making it unsuitable for WTG. Instead, one must define transversality with respect to the Weyl invariant covariant derivative. Hence, the appropriate transversality condition reads
$\tilde{\nabla}_{\mu}\xi^{\mu}=0\qquad\iff\qquad\nabla_{\mu}\xi^{\mu}=\xi^{\mu}\partial_{\mu}\ln\frac{\sqrt{-\mathfrak{g}}}{\omega}.$ (7)
Since the Lie derivative of the volume $D$-form $\bm{\omega}$ yields $\pounds_{\xi}\bm{\omega}=\bm{\omega}\tilde{\nabla}_{\mu}\xi^{\mu}$ (this result can be obtained by direct computation and also follows from the fact that $\tilde{\nabla}_{\mu}\bm{\omega}=0$), we can understand the transversality condition as defining the volume preserving transformations. Transverse diffeomorphisms transform the dynamical metric $g_{\mu\nu}$ in the usual way
@@
To find equations of motion for WTG, we vary action (5) with respect to the dynamical metric $g^{\mu\nu}$. Variation of the auxiliary Ricci scalar term yields, up to total divergence terms, the contribution $\left(\sqrt{-\mathfrak{g}}/\omega\right)^{-2/D}\omega\left(\tilde{R}_{\mu\nu}-\tilde{R}\tilde{g}_{\mu\nu}/D\right)$, where the second term comes from the variation of the factor $\left(\sqrt{-\mathfrak{g}}/\omega\right)^{-2/D}$ with respect to $g_{\mu\nu}$. Variation of the matter action similarly yields $\left(\sqrt{-\mathfrak{g}}/\omega\right)^{-2/D}\omega\left(\tilde{T}_{\mu\nu}-\tilde{T}\tilde{g}_{\mu\nu}/D\right)$, where we define the WTDiff-invariant energy-momentum tensor
@@
Dividing by the common factor $\left(\sqrt{-\mathfrak{g}}/\omega\right)^{-2/D}\omega$, we obtain the traceless, WTDiff-invariant equations of motion
$\tilde{R}_{\mu\nu}-\frac{1}{D}\tilde{R}\tilde{g}_{\mu\nu}=8\pi G\left(\tilde{T}_{\mu\nu}-\frac{1}{D}\tilde{T}\tilde{g}_{\mu\nu}\right).$ (11)
@@
$\tilde{\nabla}_{\nu}\tilde{T}_{\mu}^{\;\>\nu}=\tilde{\nabla}_{\mu}\mathcal{J},$ (12)
@@
Bianchi identities allow us to restate the traceless equations of motion (11) in a divergenceless form reminiscent of the equations of motion of GR
@@
where $\Lambda$ is an arbitrary integration constant. This represents the principal difference from general relativity, where the cosmological constant is an (arbitrary) fixed parameter in the gravitational action. The divergenceless tensor
@@
acts as the matter source for the Einstein tensor. Hence, $\tilde{T}^{\prime}_{\mu\nu}$ is the relevant tensor for coupling of matter fields to WTG, as we will see in our discussion of the EP. Let us note that, given the traceless part of the energy-momentum tensor, the corresponding divergenceless tensor $\tilde{T}^{\prime}_{\mu\nu}$ is defined uniquely up to a shift by a term of the form $C\tilde{g}_{\mu\nu}$, with $C$ being an arbitrary constant. Since this constant shift can be absorbed into the arbitrary integration constant $\Lambda$, $\tilde{T}^{\prime}_{\mu\nu}$ is completely fixed by the traceless gravitational equations of motion.
@@
The integration constant $\Lambda$ plays the role of the cosmological constant. In contrast to GR, $\Lambda$ is unrelated to any parameter of the Lagrangian and is only defined on shell, generically taking different values for various solutions of the equations of motion. In other words, $\Lambda$ represents a global degree of freedom of the theory.
These WTDiff-invariant theories belong to the so called unimodular theories of gravity. Note that fixing the gauge in WTG will bring us to unimodular gravity theory, that is not longer invariant under Weyl transformations [34, 25]. These scenarios can be useful for example in cosmology, although in other more fundamental works you face losing some information and consistency, as it would happen if you do it in electrodynamics.
@@
In the same way that there exist modified Diff-invariant theories of gravity, there exist WTDiff-invariant theories generalising WTG. Their Lagrangian can be any scalar constructed from the auxiliary metric, $\tilde{g}_{\mu\nu}$, the auxiliary Riemann tensor, $\tilde{R}^{\mu}_{\;\>\nu\rho\sigma}$, the Weyl covariant derivative $\tilde{\nabla}_{\mu}$ and some collection of matter fields, in principle non-minimally coupled. One can in fact show that for each local, WTDiff-invariant action there exists a corresponding local, Diff-invariant action which implies the same classical gravitational dynamics (except for the different behaviour of $\Lambda$) and vice versa [23]. In other words, there exist pairs of Diff and WTDiff-invariant theories of gravity that are mutually equivalent in the same sense in which GR is equivalent to WTG. As a special case, there exist WTDiff-invariant Lanczos-Lovelock theories [35, 36, 37] which are the only purely metric, WTDiff-invariant theories with second order equations of motion. They have the same classical solutions as the corresponding Diff-invariant Lanczos-Lovelock gravitational theories [37]. As we comment in section IV, Lanczos-Lovelock gravity, both Diff- and WTDiff-invariant, is also notable for obeying the strong equivalence principle.
II.1 Example: universe filled with a barotropic perfect fluid
To conclude this section, we demonstrate the key features of Weyl transverse gravity on the instructive example of a homogeneous, isotropic, spatially flat universe filled with a barotropic perfect fluid with an equation of state $p=\left(\gamma-1\right)\rho$ where $\gamma\in\left[1,2\right]$. The metric we consider reads
$\text{d}s^{2}=-\text{d}t^{2}+a^{2}\left(t\right)\text{d}x_{i}\text{d}x^{i},$ (15)
where $a\left(t\right)$ is the scale factor. The traceless part of the WTDiff-invariant energy-momentum tensor of the perfect fluid reads
$\tilde{T}_{\mu\nu}-\frac{1}{D}\tilde{T}\tilde{g}_{\mu\nu}=\left(\rho+p\right)\left(\tilde{u}_{\mu}\tilde{u}_{\nu}+\frac{1}{D}\tilde{g}_{\mu\nu}\right),$ (16)
where $\tilde{u}_{\mu}$ the (WTDiff-invariant) velocity of the fluid. The traceless equations of motion (11) yield only one independent equation (compared to two equations of general relativity)
$\dot{H}=-4\pi\left(\rho+p\right)=-4\pi\gamma\rho,$ (17)
where the overdot denotes a time derivative and $H=\dot{a}/a$ is the Hubble rate. The right hand side of this equation depends only on the combination $\rho+p$, not separately on $\rho$ and $p$ as for general relativity. The divergence condition (12) then provides the second independent equation
$\gamma\dot{\rho}+3H\rho=-\dot{\mathcal{J}}.$ (18)
If $\mathcal{J}=0$, we obtain $\rho=\rho_{0}/a^{3\gamma}$, where $\rho_{0}$ is an integration constant. Plugging this result into equation (17), multiplying it by $H$, and integrating, then gives us the Friedmann equation
$H^{2}=\frac{8\pi}{3}\rho+\frac{\Lambda}{3},$ (19)
where $\Lambda$ is an integration constant. Then, we precisely recover the dynamics of general relativity, except for the above discussed different origin of the cosmological constant $\Lambda$. If $\mathcal{J}\neq 0$, we can simply switch to the divergenceless energy-momentum tensor $\tilde{T}^{\prime}_{\mu\nu}$ given by equation (14). For our perfect fluid, it reads
$\tilde{T}^{\prime}_{\mu\nu}=\left(\rho+p\right)\tilde{u}_{\mu}\tilde{u}_{\nu}+\left(p-\mathcal{J}\right)\tilde{g}_{\mu\nu}.$ (20)
This tensor then corresponds to pressure $p-\mathcal{J}$ and energy density $\rho+\mathcal{J}$. Since the combination $\rho+p$ is unchanged, the traceless part of this energy-momentum tensor is the same as of the original $\tilde{T}_{\mu\nu}$ (see equation (16)). Therefore, on the level of the gravitational dynamics of Weyl transverse gravity, one cannot tell apart $\tilde{T}^{\prime}_{\mu\nu}$ and $\tilde{T}_{\mu\nu}$ as only the traceless part of the energy-momentum tensor is relevant. In this case, we recover the dynamics equivalent to general relativity, coupled to the divergenceless energy-momentum tensor $\tilde{T}^{\prime}_{\mu\nu}$. Although the case we discussed is very simple, the same qualitative picture holds in generic spacetimes and for arbitrary matter contents.